Two low time technos with the same issue . .

tadawsontadawson Registered User
edited December 2011 in HES Automated Lighting
I have two used Technos with the 1.3 boards that have the same issue - neither will initialize the processor at all, giving consistent but random junk on the displays.

The odd thing is that these two fixtures were taken out of service in full working condition, packed, and stored in an environmentally conditioned space for about a year, and then shipped to us. Total time on the two was stated to be about 40 hours, and considering the source, I believe that to be true - they are basically in "New in box" condition.

They were run initially on 120V/60Hz, and that is what we have been using here, and the voltage jumpers are correct.

So, the quesion is simply: What can hit two units the same way, in a year of inactivity like this?

Oh, and the 5V supplies are at 4.9 or so, and don't have the diode on them to fail - all power looks good on the logic card . . . .

Anyone have any ideas? Two with an identical issue makes me think this has got to have been seen before . . .

- Tim

Comments

  • LightGuy48LightGuy48 Registered User
    edited November 2011
    I've had this happen before, turned out the caps on the oscillator for the uProc had gone bad and the crystal was no longer oscillating.
  • tadawsontadawson Registered User
    edited November 2011
    That's certainly what this looks like. The major problem is getting to them - both boards have a big gob of clear silicone on top of the oscillator sections of the board . . . . Any suggestions on getting through that, other than careful trimming?

    Part of me is tempted to just strip the discrete oscillator, and install the X2 component - the integrated oscillator in a can . . . the only problem is that on the schematic I have, the part number is unreadable. The other caps, however, appear to be 22pF . . .

    - Tim
  • LightGuy48LightGuy48 Registered User
    edited November 2011
    You have to gently remove it, I've heard of some people using a heat gun to help soften it but I did not do this.

    I was told it's there to protect the circuit from contaminates such as fog and hazer juice which could affect the clock circuit.

    I repaired my with a low wattage soldering iron and some desoldering braid as I don't have an SMD re-work station. I then had a helper hold the new cap in place while I tack soldered it in place.
  • tadawsontadawson Registered User
    edited November 2011
    It sounds like you only replaced one of the two caps - was it C15 or C16? Or both?

    And I have SMD rework gear, it's just not wanting to do damage to the card. I have considered using the hot air nozzle on the station to soften the gunk, but if anyone had a better idea, I'm all ears!

    - Tim
  • LightGuy48LightGuy48 Registered User
    edited November 2011
    I believe I did both, it's been about 6-7 years since I did the work.
  • tadawsontadawson Registered User
    edited November 2011
    Thanks for the info . . . a dead main oscillator has been my suspect for a while, but was not willing to dig it out without confirmation of someone else having see it, which you have provided.

    More info after I get parts and can make the changes and test. . .

    - Tim
  • quinnquinn Registered User
    edited November 2011
    tadawson wrote: »
    And I have SMD rework gear, it's just not wanting to do damage to the card. I have considered using the hot air nozzle on the station to soften the gunk, but if anyone had a better idea, I'm all ears!
    these are my go-to's for removal and re-application. i've seen the stripper separate but not dissolve already-cracked soldermask, once.
  • LightGuy48LightGuy48 Registered User
    edited November 2011
    Yep, definitely been there and have the t-shirt, it started out on mine as slow to boot, but after the fixture warmed up for 5-10 minutes it would boot and run fine. Eventually it just got to the point where it would never boot.

    Replacing the two caps on mine fixed it, as I recall the crystal was fine.
  • tadawsontadawson Registered User
    edited November 2011
    These sat in storage for a year or so, and came out dead - no attempt to boot. Units are pretty much new condition, so a degraded part would appear to be the issue, and this makes sense, and is cheap to try. For quinn, my supplier shows that product as discontinued, but they have others - thanks!
  • quinnquinn Registered User
    edited November 2011
    tadawson wrote: »
    my supplier shows that product as discontinued, but they have others - thanks!
    the new MSDS shows a packaging change from 100ml to 55ml, thanks for the heads up.

    i'd definitely suggest a gel-type for getting through a thick, localized blob rather than a liquid suited for entire-board treatments.
  • tadawsontadawson Registered User
    edited November 2011
    The supplier I use has a product that is available either in a bottle, or a pen, for local removal. I was tending towards the pen, since I only need to unmask two small caps - 0606 size, if I recall (but I need to verify).

    (Glad I checked - these are 1206 form factor - HUGE for SMD, should be easy! The last I worked with were 0603 . . . that was *NOT* fun . . . )

    Also, I rough cleared the goo to check the parts, and a touch of a screwdriver to the top of C15 got the unit to boot, so I am now certain that this is the problem . . .

    So thanks again for the pointer! Parts are ordered, and these should off the bench pretty quickly!

    - Tim
  • tadawsontadawson Registered User
    edited November 2011
    Turns out that the caps alone did not do it . . . It appears that the problem is the conformal coating having degraded or something . . . The first was cap changed, and still no-go. I stripped the coating off the entire oscillator to enable me to probe the resistors, and it has been working since.

    The second unit, I decided to test that theory. I did nothing but remove the goo, and it has been working fine since . . .

    - Tim
  • tadawsontadawson Registered User
    edited December 2011
    One last update, in case it may be of help to anyone. On the first unit, things were still a bit unstable. I checked and replaced the two resistors in the oscillator as well - the 475K was reading over 1Mohm, and the 20M more like 32M. Considering these fixtures had a total time of 60 hours, and were spotless, there must have been something in the conformal that tore up these parts - not much else makes sense.

    At any rate, all working 100% now, to thanks to all that helped!

    -Tim
Sign In or Register to comment.